OpenStudy (anonymous):

solve. -11(8 - x)= x - 10(5 - x)

OpenStudy (anonymous):

-88+11x=x-50+x 9x=38 x=38/9

OpenStudy (anonymous):

ok so you have to multiply both sides by 10

OpenStudy (anonymous):

no, just the parenthesis. distributive property.

OpenStudy (anonymous):

ok thanks a lot

OpenStudy (anonymous):

anytime.

OpenStudy (anonymous):

where did 9 come from?

OpenStudy (anonymous):

divide by 9 to isolate x.

OpenStudy (anonymous):

oh ok thanks

OpenStudy (anonymous):

yup.

myininaya (myininaya):

-88+11x=x-50+10x

OpenStudy (anonymous):

hello do you have a different way of doing this problem?

OpenStudy (anonymous):

she has a different way of doing everything

myininaya (myininaya):

i was just correctiing his distrubtion